find the equation in 8-3x+5(2x+)=16
Answer by nyc_function(2741)

The answer is x = -12/7.
The question is asking you to find the value of x not an equation.
They give you the equation and you need to find x, which I found
to be x = -12/7.

8 - 3x + 5(2x + 4)= 16

8 - 3x + 10x + 20 = 16

28 + 7x = 16

7x = 16 - 28

7x = -12

x = -12/7